# Wasserstein Flow Matching: Generative modeling over families of distributions

### v1

@@ 2.3.1 Bures–Wasserstein (BW) space
A known special case of the Wasserstein space is the Bures–Wasserstein space, which consists of the submanifold of non-degenerate Gaussians parameterized by means and covariances $\{(m,\Sigma):m\in\mathbb{R}^{d},\Sigma\in\mathbb{S}^{d}_{++}\}$, endowed with the Wasserstein metric. We provide a brief exposition on the geometry of the Bures–Wasserstein space and refer the interested reader to Lambert et al., (2022) for detailed calculations and explanations, as we follow their notation conventions.
The OT map between $\mu=\mathcal{N}(m_{\mu},\Sigma_{\mu})$ and $\nu=\mathcal{N}(m_{\nu},\Sigma_{\nu})$ has a closed-form (Gelbrich,, 1990):
$T_{\star}(x)\coloneqq m_{\nu}+C^{\mu\to\nu}(x-m_{\mu})\coloneqq b+\Sigma_{\mu}^{-\tfrac{1}{2}}(\Sigma_{\mu}^{\tfrac{1}{2}}\Sigma_{\nu}\Sigma_{\mu}^{\tfrac{1}{2}})^{\tfrac{1}{2}}\Sigma_{\mu}^{-\tfrac{1}{2}}(x-m_{\mu})\,.$
As this map is affine, it is clear that the McCann interpolation between two Gaussians is always Gaussian (indeed, Gaussians undergoing affine transformations remain Gaussian). More generally, we have the succinct representation of the tangent space at a point in the Bures–Wasserstein space
@@
and the exponential and logarithmic maps between two non-degenerate Gaussians are
$(a,S)\mapsto\exp_{\mu}((a,S))\coloneqq\mathcal{N}(m_{\mu}+a,(S+I)\Sigma_{\mu}(S+I))\,,$
$\nu\mapsto\log_{\mu}(\nu)\coloneqq(m_{\nu}-m_{\mu},\Sigma_{\mu}^{-\tfrac{1}{2}}(\Sigma_{\mu}^{\tfrac{1}{2}}\Sigma_{\nu}\Sigma_{\mu}^{\tfrac{1}{2}})^{\tfrac{1}{2}}\Sigma_{\mu}^{-\tfrac{1}{2}}-I)\,,$
where the exponential map requires $S\succ-I$. We also mention that the norm on the tangent space at $\mu$ in the Bures–Wasserstein space can be written as
$\|(a,S)\|^{2}_{{\rm{BW}}(\mu)}\coloneqq\|a-m_{\mu}\|^{2}+{\rm{Tr}}(S^{2}\Sigma_{\mu})$
With the above, it is easy to compute the closed-form solutions for the mean and covariance of the McCann interpolation $\mu_{t}=(T_{t})_{\sharp}\mu=\mathcal{N}(m_{t},\Sigma_{t})$, given by
$m_{t}\coloneqq(1-t)a+tb\,,\quad\Sigma_{t}\coloneqq T_{t}AT_{t}\coloneqq((1-t)I+tC^{A\to B})A((1-t)I+tC^{A\to B})\,.$ (8)
We can relate the Euclidean and Riemannian time-derivatives of $\Sigma_{t}$ through the following manipulation (the latter of which respects the exponential and logarithmic maps above):
$\dot{\Sigma}_{t}^{\rm E}=\dot{T}_{t}AT_{t}+T_{t}A\dot{T}_{t}=\dot{T}_{t}(T_{t})^{-1}T_{t}AT_{t}+T_{t}AT_{t}(T_{t})^{-1}\dot{T}_{t}=\dot{\Sigma}_{t}^{\rm{BW}}\Sigma_{t}+\Sigma_{t}\dot{\Sigma}_{t}^{\rm{BW}}\,.$
To this end, we can draw parallels to equation 7 by writing
$\dot{m}_{t}=b-a\,,\quad\dot{\Sigma}_{t}^{\rm{BW}}=(C^{A\to B}-I)((1-t)I+tC^{A\to B})^{-1}\,.$ (9)
@@ 3 Flow matching over the Wasserstein space
3.1 Training
Let $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ denote probability measures over the Wasserstein space. [^1] This implies that $\mu\sim\mathfrak{p}_{0}$ is itself a distribution (e.g., a Gaussian or a point-cloud), not a random variable. Our goal is to learn a vector field that transports the family of measures $\mathfrak{p}_{0}$ to the family $\mathfrak{p}_{1}$. To accomplish this, we pass in the McCann interpolation $\mu_{t}$ and optimal velocity field $v_{t}$ in the Riemannian FM objective equation 2, resulting in our Wasserstein FM (WFM) objective:
$\min_{\theta}\int_{0}^{1}\!\!\iint\|f_{\theta}^{\texttt{geo}}(\mu_{t},t)-v_{t}\|_{L^{2}(\mu_{t})}^{2}\,\mathrm{d}\mathfrak{p}_{0}(\mu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)\,\mathrm{d}t\,.$ (10)
Figure 2: When the number of training examples is too few, all methods collapse on the training data, though our Riemannian instantiation of BW-FM captures the covariances perfectly. In the presence of sufficiently many samples, all methods generate Gaussians along the whole spiral, and our Riemannian BW-FM algorithm produces the most consistent samples. Other methods produce Gaussians with degenerate covariance, highlighting their inability to effectively capture the geometry of the data.
As mentioned in the introduction, our two use-cases of interest are flow matching over (1) families of point-clouds, and (2) families of Gaussian distributions. While the theory outlined in Section 2.3 explicitly requires continuous distributions to ensure all objects are well-defined, the approximation of measures by point-clouds is reasonable for our applications and can be made computationally efficient courtesy of existing open-source packages (Flamary et al.,, 2021; Cuturi et al.,, 2022). In the case of Gaussian measures, the theory as described in Section 2.3.1 holds in full force. Our training algorithm is described in Algorithm 1, and Appendix D contains precise details regarding our neural network parameterization.
Data: base $\mathfrak{p}_{0}\in\mathcal{P}(\mathcal{P}(\mathbb{R}^{d}))$, target $\mathfrak{p}_{1}\in\mathcal{P}(\mathcal{P}(\mathbb{R}^{d}))$, $\texttt{geo}\in\{\text{BW},\text{PC}\}$
Init: Parameters $\theta$ of $f_{\theta}^{\texttt{geo}}$
while not converged do
@@ 3.1 Training
             $v_{t}\leftarrow(\dot{m}_{t},\dot{\Sigma}_{t}^{\rm{BW}})$ via equation 9
@@
      $\ell(\theta)\leftarrow\left\|f_{\theta}^{\texttt{geo}}(\mu_{t},t)-v_{t}\right\|_{L^{2}(\mu_{t})}^{2}$
       $\theta\leftarrow\texttt{optimizer\_step}(\theta,\ell(\theta),\nabla_{\theta}\ell(\theta))$
Algorithm 1 Wasserstein FM Trainingd
@@
Algorithm 2 ${\rm{BW}}(\mathbb{R}^{d})$ generation
Data: Trained $f_{\theta}^{\rm{PC}}$, step size $h=1/N$
Init: $\hat{\bm{X}}_{0}=\{X_{1},\ldots,X_{n}\}\sim\mathfrak{p}_{0}$
for $k=0,\dotsc,N-1$ do
       $\hat{\bm{X}}_{(k+1)h}\leftarrow\hat{\bm{X}}_{kh}+hf_{\theta}^{\rm{PC}}(\hat{\bm{X}}_{kh},kh)$
Return: $\hat{\bm{X}}_{Nh}$
Algorithm 3 Point-cloud generation
Finally, we mention that both frameworks can be modified by training via the multisample FM algorithm (Pooladian et al., 2023a, ; Tong et al.,, 2023). In brief, the idea is to augment the training regime by pairing the source and target minibatch samples according to some prescribed matching rule (instead of independent draws from both $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$). We employ this augmentation during training, which we detail in Appendix B.
3.1.1 WFM over the Bures–Wasserstein space
@@
First suppose $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ are distributions over Gaussians, meaning that a batch of samples drawn from $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ consists of a batch of mean-covariance pairs. Here, the dynamics are straightforward: the interpolant is the McCann interpolation (recall equation 8), and the velocity field over the Bures–Wasserstein manifold is also known (see equation 9). Since $\mu_{t}$ is parameterized by $(m_{t},\Sigma_{t})$, the neural network is parameterized as $f_{\theta}^{\text{BW}}:\mathbb{R}^{d}\times\mathbb{S}^{d}_{++}\to\mathbb{R}^{d}\times\mathbb{S}^{d}$, and the norm on the tangent space simplifies the computations considerably. Our final training objective becomes
$\min_{\theta}\int_{0}^{1}\!\!\iint\|f_{\theta}^{\text{BW}}((m_{t},\Sigma_{t}),t)-(\dot{m}_{t},\dot{\Sigma}_{t}^{\rm{BW}})\|_{{\rm{BW}}(\mu_{t})}^{2}\,\mathrm{d}\mathfrak{p}_{0}(\mu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)\,\mathrm{d}t\,.$ (11)
Figure 3: Environment of cells in spatial transcriptomics data can be effectively captured via the mean and covariance of gene expression from their niche (Haviv et al., 2024b, ). Applying BW to realize environments of microglia from MERFISH data (Zhang et al.,, 2021), the composition of generated niches matches the real data along inferred cortical depth; see also Figure S2.
3.1.2 WFM over distributions of point-clouds
@@
In the case of point-clouds, we lose closed-form interpolations. However, we can hope to proceed so long as we have an approximation of the optimal transport map between the point-clouds, written $\hat{T}$. There are many works on the approximation of these maps on the basis of samples; see Hütter and Rigollet, (2021); Divol et al., (2022); Manole et al., (2021); Pooladian and Niles-Weed, (2021). We stress that the goal of our approach is to have a methodology that holds for families of point-clouds of non-uniform size.
We consider two approximations of optimal transport maps, both of which are based on entropic optimal transport Cuturi, (2013), and are computationally efficient on GPUs due to Sinkhorn’s algorithm (Sinkhorn,, 1964). One approach is to round the optimal coupling to a permutation and perform the resulting interpolation. Another approach, which allows for inhomogeneous pairs of points, is to approximate $T_{\star}^{\mu\to\nu}$ using the entropic map (Pooladian and Niles-Weed,, 2021); we provide extensive background on these objects in Appendix A. To this end, let $\bm{X}$ (resp. $\bm{Y}$) represent the locations of the point-cloud $\mu\sim\mathfrak{p}_{0}$ (resp. $\nu\sim\mathfrak{p}_{1}$), and let $\hat{T}^{\mu\to\nu}$ denote the approximation of the optimal transport map. The objective equation 10 can be approximated by
$\min_{\theta}\int_{0}^{1}\!\!\iint\sum_{i}\|[f_{\theta}^{\text{PC}}(\hat{\bm{X}}_{t},t)]_{i}-[(\hat{T}^{\mu\to\nu}(\bm{X})-\bm{X})]_{i}\|_{2}^{2}\,\mathrm{d}\mathfrak{p}_{0}(\mu)\,\mathrm{d}\mathfrak{p}_{1}(\nu)\,\mathrm{d}t\,,$ (12)
where $\hat{\bm{X}}_{t}=(1-t)\bm{X}+t\hat{T}^{\mu\to\nu}(\bm{X})$. Here, we stress that $\hat{\bm{X}}_{t}$ plays the role of a discretized McCann interpolation $\mu_{t}$.
We parameterize $f_{\theta}^{PC}$ with a transformer and Appendix D provides further details. As both the OT map and self-attention are permutation equivariant, transformers are an organic backbone for OT based models (Haviv et al., 2024a, ). Moreover, unlike other point-cloud neural models such as PVCNN (Liu et al.,, 2019), transformers do not rely on voxelization and are not hindered by the curse-of-dimensionality.
@@ 3.2 Generation
Once $f_{\theta}^{\texttt{geo}}$ is trained, we can generate new samples in a simulation-free manner as in Riemannian FM. For the Bures–Wasserstein space, we appeal to the exponential and logarithmic maps; Section 2.3.1 and Algorithm 2. We emphasize that the appropriate Riemannian updates are crucial to obtain non-degenerate final samples. To generate point-clouds, we perform a standard Euler discretization of the learned flow over the whole collection of point-clouds; see Algorithm 3.
@@ 4 Results
4.1 Flow matching between families of Gaussians
We first demonstrate our flow matching framework between measures of Gaussian distributions on a synthetic and real datasets. For comparable baselines in each scenario, we construct two simpler flow matching approaches for Gaussian generation: (1) Frobenius FM, which concatenates the mean and covariances values, and trains on $\dot{\Sigma}_{t}^{{\rm{E}}}$ with respect to the squared-Frobenius norm, and (2) BW-FM (Euclidean), which tries to match $\dot{\Sigma}^{{\rm E}}_{t}$ but still under the BW geometry (and not a concatenated geometry).
In all cases, we assess the quality of the learned flows by computing the the minimum distance between each generated Gaussian and the dataset using the (squared) $2$-Wasserstein distance. Notably, the flows generated by Frobenius FM and BW-FM (Euclidean) do not strictly adhere to the geometry of the Bures-Wasserstein manifold, requiring synthesized covariance matrices to be artificially projected onto the space of positive semi-definite (PSD) matrices via eigenvalue truncation. In contrast, the Riemannian BW-FM algorithm consistently produces valid and accurate results across all dimensions and datasets.
4.1.1 Toy Datasets
As a first test, we design a dataset of Gaussians centered on a spiral (Figure 2). When there were only few samples, only Riemannian BW-FM reconstructed the data, despite other benchmark methods following identical training regimes. On the complete 128-sample dataset, BW-FM not only reconstructs the training data, but generalizes and synthesizes novel Gaussians whose means lie on the spiral with the correct covariance profile. BW-FM shares this generalization feature with standard FM, and is able to learn the structure underlying the measure from the training data.

### v2

@@ 2.3.1 Bures–Wasserstein (BW) space
A known special case of the Wasserstein space is the Bures–Wasserstein space, which consists of the submanifold of non-degenerate Gaussians parameterized by means and covariances $\{(m,\Sigma):m\in\mathbb{R}^{d},\Sigma\in\mathbb{S}^{d}_{++}\}$, endowed with the Wasserstein metric. We provide a brief exposition on the geometry of the Bures–Wasserstein space and refer to Lambert et al. (2022) for detailed calculations and explanations, as we follow their notation conventions.
The OT map between $\mathcal{N}(m_{\mu},\Sigma_{\mu})$ and $\mathcal{N}(m_{\nu},\Sigma_{\nu})$ has a closed-form expression (Gelbrich, 1990):
$T_{\star}(x)$ $\coloneqq m_{\nu}+C^{\mu\to\nu}(x-m_{\mu})$
$\coloneqq m_{\nu}+\Sigma_{\mu}^{-\tfrac{1}{2}}(\Sigma_{\mu}^{\tfrac{1}{2}}\Sigma_{\nu}\Sigma_{\mu}^{\tfrac{1}{2}})^{\tfrac{1}{2}}\Sigma_{\mu}^{-\tfrac{1}{2}}(x-m_{\mu})\,.$
As this map is affine, it is clear that the McCann interpolation between two Gaussians is always Gaussian. More generally, we have the succinct representation of the tangent space at a point in the Bures–Wasserstein space
@@
and the tangent space norm at $\mu$ can be written as:
$\|(a,S)\|^{2}_{{\rm{BW}}(\mu)}\coloneqq\|a\|^{2}+{\rm{Tr}}(S^{2}\Sigma_{\mu})$
With the above, it is straightforward to compute the McCann interpolation $\mu_{t}=(T_{t})_{\sharp}\mu=\mathcal{N}(m_{t},\Sigma_{t})$, with
$m_{t}$ $\coloneqq(1-t)m_{\mu}+tm_{\nu}\,,$ (8)
$\Sigma_{t}$ $\coloneqq T_{t}AT_{t},\,T_{t}=(1-t)I+tC^{A\to B}\,$
We can relate the Euclidean and Riemannian time-derivatives of $\Sigma_{t}$ through the following manipulation:
$\dot{\Sigma}_{t}^{\rm E}=\dot{T}_{t}AT_{t}+T_{t}A\dot{T}_{t}=\dot{\Sigma}_{t}^{\rm{BW}}\Sigma_{t}+\Sigma_{t}\dot{\Sigma}_{t}^{\rm{BW}}\,.$ (9)
To this end, we can draw parallels to (7) by writing
$\dot{m}_{t}=m_{\nu}-m_{\mu}\,,$ (10)
$\dot{\Sigma}_{t}^{\rm{BW}}=(C^{A\to B}-I)((1-t)I+tC^{A\to B})^{-1}\,.$
@@ 3 Flow matching over the Wasserstein space
Let $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ denote probability measures over the Wasserstein space. Our goal is to learn a vector field that transports the family of measures $\mathfrak{p}_{0}$ to the family $\mathfrak{p}_{1}$. WFM learns to map source to target by regressing onto Wasserstein geodesics between samples $\mu\sim\mathfrak{p}_{0}$ and $\nu\sim\mathfrak{p}_{1}$. [^1] Crucially, we are not learning the OT map between $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$. To accomplish this, we pass in the McCann interpolation $\mu_{t}$ and optimal velocity field $v_{t}$ in the Riemannian FM objective (2), resulting in our Wasserstein FM (WFM) objective:
$\min_{\theta}\int_{0}^{1}\!\!\iint\|f_{\theta}^{\texttt{geo}}(\mu_{t},t)-v_{t}\|_{L^{2}(\mu_{t})}^{2}\,\mathrm{d}\mathfrak{p}_{0}\,\mathrm{d}\mathfrak{p}_{1}\,\mathrm{d}t\,.$ (11)
See Appendix B for a derivation of this training objective. It remains to ensure that this objective describes valid conditional probability flows between families of distributions.
First and foremost, we require two levels of regularity on the distributions of distributions $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$:
  1. 1.
    Outer continuity: The source measure $\mathfrak{p}_{0}$ satisfies smoothness conditions (as in Appendix A of Chen et al. (2024)) ensuring the conditional paths $\mathfrak{p}_{t}(\cdot|\nu)$ can be marginalized over $\mathfrak{p}_{1}$ to form a valid flow
  2. 2.
    Inner continuity: For any $\mu\sim\mathfrak{p}_{0}$ and $\nu\sim\mathfrak{p}_{1}$, there exists an optimal transport map between $\mu$ and $\nu$ making the conditional flows well-defined.
The condition of “inner continuity” is fairly mild, as this is ensured for any distribution $\mu\sim\mathfrak{p}_{0}$ with density. For Gaussian distributions, inner continuity holds naturally. For general distributions, we assume continuity but work with point-clouds as empirical realizations to approximate OT maps with statistical guarantees (see Section A.3) and computational efficiency (Flamary et al., 2021; Cuturi et al., 2022). There exists a slight gap between our theoretical results which consider classic OT and our practical implementation which uses entropy-regularized OT approximations. This aligns with other geometric methods (see Algorithm 1 by Chen & Lipman (2023)) that rely on numerical approximations when exact solutions are not tractable. The “outer continuity” condition is purely technical, and it serves the same role as in prior work. Our training algorithm is described in Algorithm 1, and Appendix E contains precise details on our neural network design.
Finally, we prove that regressing onto vector fields results in valid conditional probability flows between distributions of distributions; see Section B.1 for a proof.
Proposition 3.1.
Conditional probability paths $\mathfrak{p}_{t}(\cdot|\nu)$ generated by conditional vector fields of the form $v_{t}$ (recall (7)) satisfy $\mathfrak{p}_{1}(\cdot|\nu)=\nu$.
To prove this result, we following a general recipe by Chen & Lipman (2023) with calculations specific to the Wasserstein space. First, we prove that the Wasserstein distance is a pre-metric (see Lemma B.1). Then, we prove that the conditional vector fields $v_{t}$ are defined as a function of this pre-metric (see Lemma B.2). With these calculations, the proof of Proposition 3.1 then follows immediately from Theorem 3.1 of Chen & Lipman (2023).
Algorithm 1 Wasserstein FM Training
  Data: base $\mathfrak{p}_{0}\in\mathcal{P}(\mathcal{P}(\mathbb{R}^{d}))$, target $\mathfrak{p}_{1}\in\mathcal{P}(\mathcal{P}(\mathbb{R}^{d}))$, $\text{geo}\in\{\text{BW},\text{General}\}$
  Init: Parameters $\theta$ of $f_{\theta}^{\text{geo}}$
  repeat
@@
        $v_{t}\leftarrow(\dot{m}_{t},\dot{\Sigma}_{t}^{\text{BW}})$ via equation 9
@@
     end if
     $\ell(\theta)\leftarrow\|f_{\theta}^{\text{geo}}(\mu_{t},t)-v_{t}\|^{2}_{L^{2}(\mu_{t})}$
     $\theta\leftarrow\text{optimizer\_step}(\theta,\ell(\theta),\nabla_{\theta}\ell(\theta))$
  until Convergence
Algorithm 2 BW$(\mathbb{R}^{d})$ Generation
@@
Algorithm 3 General Distribution Generation
  Data: Trained $f_{\theta}^{\text{PC}}$, step size $h=1/N$
  Init: $\mu_{0}\sim\mathfrak{p}_{0}$, $\hat{\mathbf{X}}_{0}=\{X_{1},\ldots,X_{n}\}\sim\mu_{0}$
  for $k=0,\ldots,N-1$ do
     $\hat{\mathbf{X}}_{(k+1)h}\leftarrow\hat{\mathbf{X}}_{kh}+hf_{\theta}^{\text{PC}}(\hat{\mathbf{X}}_{kh},kh)$
  end for
  Return: $\hat{\mathbf{X}}_{Nh}$
3.1 WFM over the Bures–Wasserstein space
@@
First suppose $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ are distributions over Gaussians, meaning that a batch of samples drawn from $\mathfrak{p}_{0}$ and $\mathfrak{p}_{1}$ consists of a batch of mean-covariance pairs. Here, the dynamics are straightforward: the interpolant is the McCann interpolation, and the velocity field over the Bures–Wasserstein manifold is also known (see (8) & (10), respectively). Since $\mu_{t}$ is $(m_{t},\Sigma_{t})$, the neural network is parameterized as $f_{\theta}^{\text{BW}}:\mathbb{R}^{d}\times\mathbb{S}^{d}_{++}\to\mathbb{R}^{d}\times\mathbb{S}^{d}$, and the norm on the tangent space simplifies the computations considerably. Our final training objective becomes
$\int_{0}^{1}\!\!\iint\|f_{\theta}^{\text{BW}}((m_{t},\Sigma_{t}),t)-(\dot{m}_{t},\dot{\Sigma}_{t}^{\rm{BW}})\|_{{\rm{BW}}}^{2}\,\mathrm{d}\mathfrak{p}_{0}\,\mathrm{d}\mathfrak{p}_{1}\,\mathrm{d}t\,.$
3.2 WFM over distributions of general distributions
@@
In the case of general distributions, we lose closed-form interpolations. However, we can hope to proceed so long as we have an approximation of the optimal transport map between them, written $\hat{T}$, based on empirical samples (point-clouds) drawn from the distributions. There are many works on the approximation of these maps on the basis of samples; see Hütter & Rigollet (2021); Divol et al. (2022); Manole et al. (2021); Pooladian & Niles-Weed (2021).
While existing point-cloud generative models work with uniform-sized samples from high-resolution CAD models, many datasets provide inherently variable-sized samples. Though we assume these samples come from underlying continuous distributions, we only observe finite, discrete realizations. This natural variation in sample sizes necessitates our development of WFM to handle point-clouds of non-uniform size, making WFM applicable across both high-resolution shape modeling and naturally discrete datasets.
We approximate optimal transport maps using entropic OT (Cuturi, 2013), which is computationally efficient on GPUs via Sinkhorn’s algorithm (Sinkhorn, 1964). Let $\bm{X}$ and $\bm{Y}$ represent point-clouds described by data from $\mu\sim\mathfrak{p}_{0}$ and $\nu\sim\mathfrak{p}_{1}$ respectively, and let $\hat{T}^{\mu\to\nu}$ denote the entropic approximation of the optimal transport map (see Appendix A for full details). The objective (11) can be approximated by
$\int_{0}^{1}\!\!\iint\|f_{\theta}^{\text{PC}}(\hat{\bm{X}}_{t},t)-((\hat{T}^{\mu\to\nu}(\bm{X})-\bm{X}))\|_{2}^{2}\,\mathrm{d}\mathfrak{p}_{0}\,\mathrm{d}\mathfrak{p}_{1}\,\mathrm{d}t\,,$
where $\hat{\bm{X}}_{t}=(1-t)\bm{X}+t\hat{T}^{\mu\to\nu}(\bm{X})$ represents a discretized McCann interpolation. We parameterize $f_{\theta}^{PC}$ with a transformer, leveraging the natural alignment between the permutation equivariance of both OT maps and self-attention. Unlike point-cloud models based on voxelization, transformers maintain efficiency in high dimensions.
@@ 3.3 Generation
Once $f_{\theta}^{\texttt{geo}}$ is trained, we can generate new samples as in Riemannian FM. For the Bures–Wasserstein space, we appeal to the closed-form exponential and logarithmic maps; see Algorithm 2. We emphasize that the appropriate Riemannian updates are crucial to obtain non-degenerate final samples. For general distributions, we perform a standard Euler discretization of the learned flow on their point-cloud realizations; see Algorithm 3.
@@ 4 Results
4.1 Families of Gaussians
We first demonstrate our flow matching framework between measures of Gaussians on synthetic and real datasets. For comparable baselines in each scenario, we construct two simpler flow matching approaches for Gaussian generation: (1) Frobenius FM, which concatenates the mean and covariances, and trains on $\dot{\Sigma}_{t}^{{\rm{E}}}$ (equation (9)) with respect to the squared-Frobenius norm, and (2) BW-FM (Euclidean), which tries to match $\dot{\Sigma}^{{\rm E}}_{t}$ but still under the BW geometry.